OpenStudy (anonymous):

@sammixboo what mixed number falls in between 17/5 and 7/2

sammixboo (sammixboo):

Okay first we need to make these improper fractions into mixed fractions... \(\sf \large \color{magenta}{\frac{17}{5}}\) ---- \(\sf \large \color{magenta}{3\frac{2}{5}}\) \(\sf \large \color{magenta}{\frac{7}{2}}\) --- \(\sf \large \color{magenta}{3\frac{1}{2}}\)

sammixboo (sammixboo):

Then we need to make the mixed fractions into twenties. \(\sf \large \color{magenta}{3\frac{2 x 4}{5 x 4}}\) --- \(\sf \large \color{magenta}{3\frac{8}{20}}\) \(\sf \large \color{magenta}{3\frac{1 x 10}{2 x 10}}\) ---- \(\sf \large \color{magenta}{3\frac{10}{20}}\)

sammixboo (sammixboo):

Now tell me what fraction is between \(\sf \large \color{magenta}{3\frac{8}{20}}\) and \(\sf \large \color{magenta}{3\frac{10}{20}}\)